# mO shizhe[#1]

### v1

@@ Abstract
Retrieval-augmented generation (RAG) has become a main technique for alleviating hallucinations in large language models (LLMs). Despite the integration of RAG, LLMs may still present unsupported or contradictory claims to the retrieved contents. In order to develop effective hallucination prevention strategies under RAG, it is important to create benchmark datasets that can measure the extent of hallucination. This paper presents RAGTruth, a corpus tailored for analyzing word-level hallucinations in various domains and tasks within the standard RAG frameworks for LLM applications. RAGTruth comprises nearly 18,000 naturally generated responses from diverse LLMs using RAG. These responses have undergone meticulous manual annotations at both the individual cases and word levels, incorporating evaluations of hallucination intensity. We not only benchmark hallucination frequencies across different LLMs, but also critically assess the effectiveness of several existing hallucination detection methodologies. Furthermore, we show that using a high-quality dataset such as RAGTruth, it is possible to finetune a relatively small LLM and achieve a competitive level of performance in hallucination detection when compared to the existing prompt-based approaches using state-of-the-art large language models such as GPT-4.111The RAGTruth dataset is available at https://github.com/ParticleMedia/RAGTruth
@@ 1 Introduction
To reduce hallucination, various methods have been developed that can be applied at different stages of LLM lifecycle, including pre-training Brown et al. (2020), supervised fine-tuning Zhou et al. (2023); Zhang et al. (2023a), RLHF Ouyang et al. (2022); Lin et al. (2022), and inference Dhuliawala et al. (2023); Gao et al. (2023). In terms of detection, methods are developed by examining the model’s intrinsic state Guo et al. (2017), comparing it with external data and tools Chern et al. (2023), or leveraging the LLM’s inherent powerful capabilities for self-checking Agrawal et al. (2023); Manakul et al. (2023). RAG is extensively used to supply LLMs with updated, relevant knowledge, significantly mitigating hallucination. Nevertheless, even with RAG and other enhancements, LLMs still produce statements that are either unfounded or contradict the information provided in the retrieved references Shuster et al. (2021).
Despite the growing awareness of the hallucination phenomenon, the understanding of hallucination in LLMs is still in its early stages. One key challenge is the lack of high-quality, large-scale datasets specifically designed for hallucination detection. This issue is particularly acute in RAG settings. Due to the relatively low hallucination ratio, a substantial increase in annotation resources is needed. Existing datasets for LLM hallucination detection are predominantly synthesized Li et al. (2023). For instance, in Liu and Liu (2023); Longpre et al. (2021) prompts conflicting with conventional knowledge are purposely generated to trigger hallucinations. While these approaches are efficient at generating hallucinations, the resulting artificial hallucinations can substantially differ from those that naturally occur. In Chen et al. (2023); Hu et al. (2023), hallucination datasets are developed by manual annotations of naturally produced LLM responses. However, these datasets are of limited size and are not specifically focused on the RAG scenario.
@@
    We present a simple baseline method of fine-tuning LLM for hallucination detection. It is shown that by fine-tuning the Llama-2-13B model on the RAGTruth training data, we can achieve results competitive to the existing prompt-based approaches using GPT-4. This shows the potential of developing better hallucination detection methods using RAGTruth.
@@ 2.1 Hallucination of Large Language Models
In Zhang et al. (2023b), the hallucination of large language models (LLMs) is characterized into three categories: 1) input-conflicting hallucination 2) context-conflicting hallucination 3) fact-conflicting hallucination. Existing research has demonstrated that incorporating up-to-date, relevant knowledge in the prompt can effectively reduce fact-conflicting hallucination Vu et al. (2023); Lewis et al. (2021). This approach, referred to as Retrieval-Augmented Generation (RAG), is widely used in real-world LLM applications. For instance, Google Bard  [^2] https://bard.google.com and Microsoft BingChat  [^3] https://www.bing.com have implemented this technique. Consequently, the detection and mitigation of input-conflicting and context-conflicting hallucinations within the RAG framework are becoming increasingly important.
@@ 2.2 Hallucination Detection Methods
Researchers have been exploring various methods to enhance the reliability of LLMs by detecting hallucinations. In Azaria and Mitchell (2023); Xiao and Wang (2021); Malinin and Gales (2021), intrinsic model uncertainty metrics such as token-level probability and entropy are used to detect hallucinations. When direct access to output uncertainty is not feasible, as in the case with limited APIs like GPT-4, an alternative approach involves employing a fully accessible LLM as a proxy Manakul et al. (2023). In Falke et al. (2019); Barrantes et al. (2020), natural language inference modules are adapted to check the information consistency between the articles and their summaries, and it has been shown that external knowledge is helpful for detecting factual hallucinations. Guo et al. (2022); Mallen et al. (2022). Additionally, methods that harness the inherent capabilities of LLMs have been proposed, such as verbalize-based and consistency-based methods Xiong et al. (2023); Manakul et al. (2023). These techniques aim to detect hallucinations without relying on internal states or external data and tools.
Figure 1: Data gathering pipeline. Taking a data-to-text task as an example, our data gathering pipeline includes 2 steps: 1) response generation. We generated responses with multiple LLMs and natural prompts. 2) human annotation. Human labeler annotated hallucinated spans in LLM responses.
2.3 Hallucination Evaluation Datasets
Constructing a hallucination dataset can be a labor-intensive task, as it necessitates the manual identification of ungrounded or falsely claimed statements in the output of LLMs. Such a dataset can be particularly valuable in the context of RAG scenarios, owing to its significantly lower incidence of hallucinations.
Hallucinate datasets largely fall into two categories: synthetic dataset and manually-crafted dataset. Synthetic datasets typically involve purposely triggering Language Model to produce spurious responses Li et al. (2023) or simply inserting knowledge conflicts into the output Longpre et al. (2021). While these methods are effective in generating a substantial volume of hallucinated responses, these responses may not follow the same distribution as the naturally generated responses. On the other hand, manually annotated datasets offer a closer representation of the LLMs’ true generative capabilities Vu et al. (2023); Lewis et al. (2021). However, the considerable labor and expense required for annotation lead to datasets of limited size and length and inconsistent quality. Specifically, there is a lack of manually curated datasets of substantial size and quality to facilitate hallucination detection in RAG scenarios.
@@ 3.1 Response Generation
We selected three widely recognized RAG tasks for response generation: Summarization, Question Answering, and Data-to-Text generation.
For the summarization task, we randomly selected documents from the training set of the well-known CNN/Daily Mail dataset See et al. (2017) as well as recent news articles from a prestigious news platform. LLMs were prompted to generate summaries for each of the source documents.
For the task of question answering, we conducted a random sampling from the training set of MS MACRO Nguyen et al. (2016). To reduce the annotation cost, only 3 retrieved passages for each question were preserved. Then we prompted LLMs to generate answers for each question solely based on the retrieved passages.
For the data-to-text generation task, we randomly sampled businesses in the restaurant and nightlife categories from the Yelp Open Dataset Yelp (2021). This dataset served as our structured data source. To streamline the annotation process, we focused only on the following business information fields: BusinessParking, RestaurantsReservations, OutdoorSeating, WiFi, RestaurantsTakeOut, RestaurantsGoodForGroups, Music, and Ambience. We then prompted LLMs to generate an objective text introduction for each business.
passage 2:…
passage 3:…
Question: how to prepare to get an ultrasound?
Response: Based on the given passages, here’s how to prepare for an ultrasound:* For a gender reveal ultrasound, wait until at least 17 weeks of pregnancy to ensure accurate results. * For 3D ultrasounds, schedule the appointment between 20 and 32 weeks of pregnancy for the best pictures
Hallucination or Not: Yes
Span: between 20 and 32 weeks of pregnancy for the best pictures
Type: Evident Conflict
Table 1: An example of our annotated data from Question Answering task. It contains reference, question, model generated response with hallucination label, annotated hallucination span and corresponding reason.
@@ 3.2 Human Annotation
Identifying AI-generated hallucinations is a challenging task. It requires a strong capacity for critical thinking to understand the logical flow of various texts, along with meticulous attention to detail for spotting subtle inaccuracies and inconsistencies. Moreover, a certain level of media literacy and knowledge of current affairs is crucial to grasp the subjects discussed in news-related sample data. Therefore, we chose annotators who are proficient in English and possess a bachelor’s degree in relevant fields to ensure the accuracy and reliability of the annotation results.
The annotators are invited to perform annotation tasks using Label Studio Tkachenko et al. (2020-2022). Each labeling task is presented within one page, comprising the following components: 1) The original input text provided to the AI models; 2) A set of 6 responses, generated by different AI models.
Their task was to annotate the specific spans of the generated text that contains hallucinated information and categorize them into four types, based on the criteria outlined below. For examples of each type, please refer to Appendix A.
Evident Conflict:
for when generative content presents direct contraction or opposition to the original input. These conflicts are easily verifiable without extensive context, often involving clear factual errors, misspelled names, incorrect numbers, etc.
Subtle Conflict:
for when generative content presents a departure or divergence from the original input, altering the intended contextual meaning. These conflicts often involve substitution of terms that carry different implications or severity, requiring deeper understanding of their contextual applications.
Evident Introduction of Baseless Information:
for when generated content includes information not substantiated in the original input. It involves the creation of hypothetical, fabricated, or hallucinatory details lacking evidence or support.
Subtle Introduction of Baseless Information:
is when generated content extends beyond the original input by incorporating inferred details, insights, or sentiments. This additional information lacks verifiability and might include subjective assumptions or commonly observed norms rather than explicit facts.
@@ 3.3 Annotations for Adaptive Evaluation
Hallucination annotation is not always straightforward. In contentious cases, additional annotations are provided to accurately reflect these situations. This approach enables users to adopt various evaluation strategies tailored to their specific application circumstances.
Incorrectly Refusing to Answer
In the question-answering task, the model is required to decline to answer when the context lacks relevant information. However, there are instances where the model mistakenly refuses to answer despite the presence of the relevant context. These refusal responses are not labeled as hallucinations but are instead marked as incorrect_refusal.
@@
Summarization(CNN/DM) 628 3768 648 1749 124 632 1230 32.6% 1570
Summarization(Recent News) 316 1896 369 481 89 240 540 28.5% 613
Question Answering 992 5952 243 509 119 381 1844 31.0% 3155
Data-to-text 1037 6222 354 1253 159 369 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
Overall 2973 17838 380 1749 131 632 7918 44.4% 14781
Table 2: The basic statistics of RAGTruth. $\dagger$: The data presented in the table above include hallucination spans caused by null values in JSON. If these spans are removed, the hallucination response number, hallucination response rate, and hallucination span number of date-to-text tasks are 3898, 62.6%, and 7797, respectively.
@@
In the data-to-text task, certain fields sometimes are with null values. We observed that in the generated results, null is often interpreted as false by some models. Since the more common expressions for negation in our dataset are the boolean value False or the text No, we labeled these instances as hallucinations (evident introduction of baseless info), and provided a special span-level annotation named due_to_null. Users can independently choose whether to consider these spans as hallucinations. Additionally, we have included an option in the evaluation script that allows users to decide whether to include these spans in the evaluation.
@@ 4.1 Basic Statistics
We presented detailed statistics of RAGTruth in Table 2. Compared to existing datasets for hallucination detection Cao et al. (2023); Kamoi et al. (2023), the RAGTruth dataset is considerably large in scale. The corpus contains a total of 2,973 instances of data, which include 944 instances for summarization, 992 instances for question answering, and 1,037 instances for data-to-text generation. Each instance comprises responses from 6 different models. As shown in Table 2, the RAGTruth dataset also features longer prompt and response lengths than existing datasets for hallucination detection Wang et al. (2020).
@@ 4.2 Hallucination Statistics
As shown in Figure 2, the generation of information baseless in the context was significantly more prevalent than the generation of information conflicting with the context. Within the two major categories of baseless info and conflict, the more severe hallucinations, namely Evident baseless info and Evident conflict, respectively, account for a significant portion. This observation highlights the importance and challenges of hallucination mitigation, even in RAG settings.
@@
As shown in Table 2, across the three tasks, the data-to-text task exhibited the highest frequency of hallucinations in its responses. When processing Yelp business data, the models often generated information not present in the input, influenced by stereotypes inherent in the business data. Furthermore, inconsistent handling of JSON format data also contributed to a significant number of hallucinations. In contrast, the hallucination rate for the summarization task was relatively lower. Interestingly, the models did not show a higher rate of hallucinations for recent news compared to outdated news. This could be attributed to the shorter context length in the recent news subtask compared to the CNN/DM subtask.
Model Summarization Question Answering Data-to-text Overall
@@
GPT-3.5-turbo-0613 64 74 0.07 81 97 0.13 283 397 0.18 428 568
GPT-4-0613 78 85 0.08 48 57 0.06 293 359 0.27 419 501
Llama-2-7B-chat 451 542 0.61 542 1082 0.62 899 1812 1.30 1892 3436
Llama-2-13B-chat 333 395 0.47 426 706 0.51 988 2840 1.54 1747 3941
${}^{\dagger}$ 220 254 0.27 354 583 0.44 873 1861 1.17 1447 2698
Mistral-7B-Instruct 624 842 0.86 393 630 0.61 968 2174 1.53 1985 3646
@@
CLB Summarization Data2text QA
1 $0.31_{(102,317]}$ $1.55_{(117,237]}$ $0.55_{(79,163]}$
2 $0.33_{(317,387]}$ $1.49_{(237,290]}$ $0.45_{(163,196]}$
3 $0.38_{(387,481]}$ $1.54_{(290,355]}$ $0.53_{(196,266]}$
4 $0.41_{(481,810]}$ $1.49_{(355,448]}$ $0.58_{(266,327]}$
5 $0.50_{(810,1749]}$ $1.53_{(448,1253]}$ $0.54_{(327,509]}$
RLB Summarization Data2text QA
1 $0.36_{(7,74]}$ $1.24_{(35,120]}$ $0.14_{(1,70]}$
2 $0.34_{(74,92]}$ $1.21_{(120,136]}$ $0.26_{(70,98]}$
3 $0.33_{(92,110]}$ $1.63_{(136,161]}$ $0.38_{(98,127]}$
4 $0.48_{(110,144]}$ $1.85_{(161,204]}$ $0.67_{(127,166]}$
5 $0.42_{(144,632]}$ $1.67_{(204,369]}$ $1.21_{(166,381]}$
@@
To more clearly compare the hallucination rate of different models, we calculated the hallucination density for each model across three tasks. Hallucination density is defined as the average number of hallucination spans per hundred words in a response. In the Llama2 series, a clear negative correlation was observed between the model scale and the hallucination density. Despite its strong performance in various benchmarks and leaderboards Zheng et al. (2023), the Mistral-7B-Instruct model recorded the highest hallucination density among the evaluated models.
@@
We divided the data for each type of task into five equal buckets based on context/response length and calculated the average number of hallucination spans per response within each bucket. As shown in Table 4, with a few exceptions, there is a clear overall trend of an increase in the average number of hallucinations as the context length or response length grows.
@@
In Figure 3, we present an analysis of the distribution of hallucinatory spans within responses. Generally, hallucinations are more likely to occur towards the end of responses, a pattern that is particularly evident in question-answering and summarization tasks. Compared to other tasks, the data-to-text task has a relatively higher occurrence of hallucinations in the first half, as indicated by the two bright areas on the heatmap.
Figure 3: Heatmaps of normalized hallucination occurrence positions. The numerical values on the colorbar represent the proportion of hallucinations occurring at a specific location relative to the total number of hallucinations in that task.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) has become a main technique for alleviating hallucinations in large language models (LLMs). Despite the integration of RAG, LLMs may still present unsupported or contradictory claims to the retrieved contents. In order to develop effective hallucination prevention strategies under RAG, it is important to create benchmark datasets that can measure the extent of hallucination. This paper presents RAGTruth, a corpus tailored for analyzing word-level hallucinations in various domains and tasks within the standard RAG frameworks for LLM applications. RAGTruth comprises nearly 18,000 naturally generated responses from diverse LLMs using RAG. These responses have undergone meticulous manual annotations at both the individual case and word levels, incorporating evaluations of hallucination intensity. We not only benchmark hallucination frequencies across different LLMs, but also critically assess the effectiveness of several existing hallucination detection methodologies. We show that using a high-quality dataset such as RAGTruth, it is possible to finetune a relatively small LLM and achieve a competitive hallucination detection performance when compared to the existing prompt-based approaches using state-of-the-art LLMs such as GPT-4. Furthermore, the finetuned model can effectively mitigate hallucination in LLM responses.
@@ 1 Introduction
Question: how to prepare to get an ultrasound?
Response: Based on the given passages, here’s how to prepare for an ultrasound:
For a gender reveal ultrasound, wait until at least 17 weeks of pregnancy to ensure accurate results.
* For 3D ultrasounds, schedule the appointment between 20 and 32 weeks of pregnancy for the best pictures
Annotation
Span: between 20 and 32 weeks of pregnancy for the best pictures
Type: Evident Conflict
Reason: Original: "the best pictures are between 24 and 30 weeks", Generative: "between 20 and 32 weeks of pregnancy for the best pictures"
Table 1: An example of RAGTruth data from the question answering task. It contains context, response generated by LLM with and span-level annotation.
To reduce hallucination, various methods have been developed that can be applied at different stages of LLM lifecycle, including pre-training Brown et al. (2020), supervised fine-tuning Zhou et al. (2023); Zhang et al. (2023a), RLHF Ouyang et al. (2022); Lin et al. (2022), and inference Dhuliawala et al. (2023); Gao et al. (2023). In terms of detection, methods are developed by examining the model’s intrinsic state Guo et al. (2017), comparing it with external data and tools Chern et al. (2023), or leveraging the LLM’s inherent powerful capabilities for self-checking Agrawal et al. (2023); Manakul et al. (2023). Retrieval-augmented generation (RAG) is extensively used to supply LLMs with updated, relevant knowledge, significantly mitigating hallucination Varshney et al. (2023). Nevertheless, even with RAG and other enhancements, LLMs still produce statements that are either unfounded or contradict the information provided in the retrieved references Shuster et al. (2021).
Despite the growing awareness of the hallucination phenomenon, the understanding of hallucination in LLMs is still in its early stages. One key challenge is the lack of high-quality, large-scale datasets specifically designed for hallucination detection. This issue is particularly acute in RAG settings. Due to the relatively low hallucination ratio, a substantial increase in annotation resources is needed. Existing datasets for LLM hallucination detection are predominantly synthesized Li et al. (2023). For instance, in Liu and Liu (2023); Longpre et al. (2021), prompts conflicting with conventional knowledge are purposely generated to trigger hallucinations. While these approaches are efficient at generating hallucinations, the resulting artificial hallucinations can substantially differ from those that naturally occur. In Chen et al. (2023); Hu et al. (2023), hallucination datasets are developed by manual annotations of naturally produced LLM responses. However, these datasets are of limited size and are not specifically focused on the RAG scenario.
@@
    We present a baseline method of fine-tuning LLM for hallucination detection. It is shown that by fine-tuning the Llama-2-13B model on the RAGTruth training data, we can achieve results competitive to the existing prompt-based approaches using GPT-4. This shows the potential of developing better hallucination detection methods using RAGTruth.
@@ 2.1 Hallucination of Large Language Models
Though hallucination in traditional natural language generation (NLG) contexts has been widely studiedJi et al. (2023), comprehending and tackling this problem in the context of LLMs presents distinct challengesZhang et al. (2023b). Existing research has demonstrated that incorporating up-to-date, relevant knowledge in the prompt can effectively reduce fact-conflicting hallucination Vu et al. (2023); Lewis et al. (2021). This approach, referred to as Retrieval-Augmented Generation (RAG), is widely used in real-world LLM applications. For instance, Google Bard  [^1] https://bard.google.com and Microsoft BingChat  [^2] https://www.bing.com have implemented this technique.
2.2 Hallucination Evaluation Datasets
Extensive research has focused on hallucination benchmarks within conventional Natural Language Generation settings Dziri et al. (2022); Zhong et al. (2021); Durmus et al. (2020); Lin et al. (2022). With the rise of LLMs, the detection of hallucinations has become increasingly challenging, necessitating the development of high-quality datasets for LLM evaluation Chen and Shu (2023). Contributions in this domain include HaluEval Li et al. (2023), which introduced datasets encompassing both synthetically and naturally generated LLM responses, and FELM Chen et al. (2023), which concentrated on naturally generated LLM responses across multiple domain tasks. RefChecker Hu et al. (2023), a distinctive approach, breaks down claims in LLM responses into triples and utilizes human annotation to assess the truthfulness of facts. Notably, these works primarily focus on annotating factual hallucinations in LLM responses. Distinguishing from previous research, our work centers on the evaluation of LLMs within RAG settings.
Figure 1: Data gathering pipeline. Taking a data-to-text writing task as an example, our data gathering pipeline includes 2 steps: 1) response generation. We generated responses with multiple LLMs and natural prompts. 2) human annotation. Human labeler annotated hallucinated spans in LLM responses.
@@ 2.3 Hallucination Detection Methods
Researchers have been exploring various methods to enhance the reliability of LLMs by detecting hallucinations. In Azaria and Mitchell (2023); Xiao and Wang (2021); Malinin and Gales (2021), intrinsic model uncertainty metrics such as token-level probability and entropy are used to detect hallucinations. When direct access to output uncertainty is not feasible, as in the case with limited APIs like GPT-4, an alternative approach involves employing a fully accessible LLM as a proxy Manakul et al. (2023). In Falke et al. (2019); Barrantes et al. (2020), natural language inference modules are adapted to check the information consistency between the articles and their summaries, and it has been shown that external knowledge is helpful for detecting factual hallucinations. Guo et al. (2022); Mallen et al. (2022). Additionally, methods that leverage the inherent capabilities of LLMs have been proposed for self-checking, such as verbalization-based and consistency-based methods Xiong et al. (2023); Manakul et al. (2023). These techniques aim to detect hallucinations without relying on internal states or external data and tools.
@@ 3 Construction Process of RAGTruth
3.1 Hallucination Taxonomy
Different from open-end generation, under RAG setting, the prompt contains rich context information, and the model is generally required to generate text based on the provided context. The detection and mitigation of inconsistencies between retrieved information and responses emerge as significant sources of hallucination.
As outlined below, we categorize the hallucination in the RAG setting into four types. For concrete examples of each type, please refer to Appendix A.
Evident Conflict:
for when generative content presents direct contraction or opposition to the provided information. These conflicts are easily verifiable without extensive context, often involving clear factual errors, misspelled names, incorrect numbers, etc.
Subtle Conflict:
for when generative content presents a departure or divergence from the provided information, altering the intended contextual meaning. These conflicts often involve substitution of terms that carry different implications or severity, requiring a deeper understanding of their contextual applications.
Evident Introduction of Baseless Information:
for when generated content includes information not substantiated in the provided information. It involves the creation of hypothetical, fabricated, or hallucinatory details lacking evidence or support.
Subtle Introduction of Baseless Information:
is when generated content extends beyond the provided information by incorporating inferred details, insights, or sentiments. This additional information lacks verifiability and might include subjective assumptions or commonly observed norms rather than explicit facts.
@@ 3.2 Response Generation
We selected three widely recognized generation tasks with RAG settings for response generation: Question Answering, Data-to-text Writing, and News Summarization.
For the task of question answering, we conducted a random sampling from the training set of MS MARCO Nguyen et al. (2016). To reduce the difficulty of annotation, we selected only those questions related to daily life, and preserved only three retrieved passages for each question. Then we prompted LLMs to generate answers for each question solely based on the retrieved passages.
For the data-to-text writing task, we prompted LLMs to generate an objective overview for a randomly sampled business in the restaurant and nightlife categories from the Yelp Open Dataset Yelp (2021). In this dataset, information pertaining to a business is represented using structured data. To streamline the annotation process, we focused only on the following business information fields: BusinessParking, RestaurantsReservations, OutdoorSeating, WiFi, RestaurantsTakeOut, RestaurantsGoodForGroups, Music, and Ambience. In addition to the structured data, we have also included up to three business-related user reviews to enrich the context information. In the prompt, these information is represented in JSON format.
For the news summarization task, we randomly selected documents from the training set of the well-known CNN/Daily Mail dataset See et al. (2017) as well as recent news articles from a prestigious news platform. LLMs were prompted to generate a summary for each of the source news.
@@ 3.3 Human Annotation
Identifying AI-generated hallucinations is a challenging task. It requires a strong capacity for critical thinking to understand the logical flow of various texts, along with meticulous attention to detail for spotting subtle inaccuracies and inconsistencies. Moreover, a certain level of media literacy and knowledge of current affairs is crucial to grasp the subjects discussed in news-related sample data. Therefore, we chose annotators who are proficient in English and possess a bachelor’s degree in English, Communications, or relevant fields to ensure the accuracy and reliability of the annotation results. We recruited annotators from a professional vendor and paid them at a rate of \$25 per hour per individual.
The annotators are invited to perform annotation tasks using Label Studio Tkachenko et al. (2020-2022). Each labeling task is presented within one page, comprising the following components: 1) the context provided to the AI models; 2) a set of 6 responses, generated by different AI models. Our annotation interface is available in Appendix C.
Their task was to annotate the specific spans of the generated text that contains hallucinated information and categorize them into the four types. To ensure the quality of the annotations, each response is independently labeled by two annotators. The consistency rate of two annotators was 91.8% at the response level and 78.8% at the span level. In cases where there is a considerable difference between the two annotations, a third review is undertaken.
@@ 3.4 Annotations for Adaptive Evaluation
In different contexts, the definition and criteria for hallucination vary, and the annotation of hallucination is not always straightforward. In contentious cases, additional annotations are provided to accurately reflect these situations. This approach enables users to adopt various evaluation strategies tailored to their specific application circumstances. Please refer to Appendix C for more statistical information about these annotations.
Implicit Truth
The extensive world knowledge and ability of LLMs is a significant advantage in open-ended generation scenarios. But in the context of this paper, which focuses on the relatively strict RAG scenarios, we have labeled information that is not mentioned in the reference but may be truthful as hallucinations. For instance, mentioning a local officer’s name not present in the reference or claiming that a restaurant accepts credit card payments without any basis.
The decision is based on the observation that LLMs have a relatively high chance of making errors when generating detailed facts, partly because their embedded knowledge can be outdated. Therefore, RAG applications usually instruct LLMs not to generate factual content without the support of references. Besides, we provided an additional span-level annotation named implicit_true for these spans to accommodate different application needs.
@@
Question Answering 989 5934 243 509 119 381 1724 29.1% 2927
Data-to-text Writing 1033 6198 354 1253 159 369 4254 68.6% 9290
Summarization(CNN/DM) 628 3768 648 1749 124 632 1165 30.9% 1474
Summarization(Recent News) 315 1890 369 481 89 240 521 27.6% 598
Overall 2965 17790 381 1749 131 632 7664 43.1% 14289
Table 2: The basic statistics of RAGTruth. Here "Resp." stands for "Response".
@@
In the data-to-text writing task, certain fields sometimes are with null values. We observed that in the generated results, null is often interpreted as false by some models. Since the more common expressions for negation in our dataset are the boolean value False or the text No, we labeled these instances as hallucinations (evident introduction of baseless info) and provided a special span-level annotation named due_to_null for these spans. In the subsequent hallucination detection experiments, our prompts will be aligned with this standard.
@@ 4.1 Basic Statistics
We presented detailed statistics of RAGTruth in Table 2. Compared to existing datasets for hallucination detection Cao et al. (2023); Kamoi et al. (2023), the RAGTruth dataset is considerably large in scale. The corpus contains a total of 2,965 instances of data, which include 989 instances for question answering, 1,033 instances for date-to-text writing, and 943 instances for news summarization. Each instance comprises responses from 6 different models. As shown in Table 2, the RAGTruth dataset also features longer prompt and response lengths than existing datasets for hallucination detection Wang et al. (2020).
@@ 4.2 Hallucination Statistics
As shown in Figure 2, the generation of information baseless in the context was significantly more prevalent than the generation of information conflicting with the context, especially for the question answering tasks. Within the two major categories of baseless info and conflict, the more severe hallucinations, namely Evident baseless info and Evident conflict, respectively, account for a significant portion. This observation highlights the importance and challenges of LLMs hallucination mitigation, even in RAG settings.
@@
As shown in Table 2, across the three tasks, the date-to-text writing task exhibited the highest frequency of hallucinations in its responses. Inconsistent handling of JSON format data, especially time and attributes, contributed to a significant number of hallucinations in this task. Interestingly, the models did not show a higher rate of hallucinations for recent news compared to outdated news. This could be attributed to the shorter context length in the recent news subtask compared to the CNN/DM subtask.
Model Question Answering Data-to-text Writing Summarization Overall
@@
GPT-3.5-turbo-0613 75 89 0.12 272 384 0.18 54 60 0.05 401 533
GPT-4-0613 48 51 0.06 290 354 0.27 74 80 0.08 406 485
Llama-2-7B-chat 510 1010 0.59 888 1775 1.27 434 517 0.58 1832 3302
Llama-2-13B-chat 399 654 0.48 983 2803 1.53 295 342 0.41 1677 3799
Llama-2-70B-chat 320 529 0.40 863 1834 1.15 212 245 0.26 1395 2608
Mistral-7B-Instruct 378 594 0.59 958 2140 1.51 617 828 0.86 1953 3562
@@
CLB Summarization D2T Writing QA
1 $0.29_{(176,368]}$ $1.51_{(178,273]}$ $0.50_{(131,187]}$
2 $0.36_{(368,587]}$ $1.48_{(273,378]}$ $0.51_{(187,288]}$
3 $0.44_{(587,1422]}$ $1.49_{(378,731]}$ $0.49_{(288,400]}$
RLB Summarization D2T Writing QA
1 $0.34_{(44,87]}$ $1.20_{(93,131]}$ $0.21_{(19,93]}$
2 $0.32_{(87,119]}$ $1.59_{(131,175]}$ $0.37_{(93,138]}$
3 $0.44_{(119,245]}$ $1.69_{(175,258]}$ $0.87_{(138,257]}$
@@
To more clearly compare the hallucination rate of different models, we calculated the hallucination density for each model across three tasks. Hallucination density is defined as the average number of hallucination spans per hundred words in the responses. In the Llama2 series, a clear negative correlation was observed between the model scale and hallucination density, aside from the data-to-text writing tasks. Despite its strong performance in various benchmarks and leaderboards Zheng et al. (2023), the Mistral-7B-Instruct model generated the highest number of responses containing hallucinations.
@@
After removing the top and bottom 5% of outliers, we partitioned the data for each task type into three equal-sized groups according to the length of the context/response. We then computed the average number of hallucinated spans per response within each group. As shown in Table 4, there is a clear overall trend of an increase in the average number of hallucinations as the response length grows. Only the average number of hallucinations in news summarization tasks significantly increases with the length of the context. This may be because the contexts in the other two tasks are more structured, and an increase in length does not significantly raise the difficulty of understanding the content.
@@
In Figure 3, we present the heatmap of the hallucination occurrence positions. Hallucinations are significantly more likely to occur towards the end of responses in question-answering and news summarization tasks. Compared to other tasks, the data-to-text writing task has a relatively higher occurrence of hallucinations in the first half. In that bright area, hallucinations concerning business attributes frequently occur.
Figure 3: Heatmaps of normalized hallucination occurrence positions. The probability of hallucinations occurring is higher in brighter areas.